OpenStudy (anonymous):

Need help: ln x + ln (x-6) = 2

OpenStudy (anonymous):

Can you please explain how you do it etc, because I would rather understand it than just get a straight answer

OpenStudy (anonymous):

I haven't taken a math class in a few months, so I forgot some of the material

OpenStudy (anonymous):

We have: \[ \ln x+\ln(x-6)=2\implies\\ \ln x(x-6)=2\implies\\ e^{\ln x(x-6)}=e^2 \](Considering the sums of logs and their definition): \[ x(x-6)=e^2\implies\\ x^2-6x=e^2\implies\\ x^2-6x-e^2=\;0 \]Using quadratic formula we get: \[ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{6\pm\sqrt{36-4e^2}}{2}=3\pm\sqrt{9+e^2} \]The solution cannot be negative, since \(\ln x\) is not defined as real for negative \(x\), so: \[ x=3+\sqrt{9+e^2} \]Et voilá.

OpenStudy (anonymous):

Ah okay. I got to the x^(2) - 6x - e^(2) = 0 part but I didnt know what to do from there. Thank you so much :)

OpenStudy (anonymous):

Sure thing.